OpenStudy (anonymous):

Find 2r2 + 3s3 - r2 + 4t2 - r2 if r = -2, s = -3, and t = 5.

OpenStudy (bibby):

simplify each term on its own. 2(r^2)=2((-2)^2)=2(4)=8 Continue. Try 3s^3